OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (1,4) and (5,12)

OpenStudy (tkhunny):

Can you calculate the slope?

OpenStudy (anonymous):

It doesn't say

OpenStudy (anonymous):

2

OpenStudy (anonymous):

do you need the answer of the slope or just the equation? 12-4 over 5-1 - answer is 2

OpenStudy (anonymous):

just the equation

OpenStudy (tkhunny):

You do not need just the equation. You do not need just the slope. You need to understand what you are doing. The very first task is to calculate the slope. You must have this skill or you will fail this course. Do you know how to do it? (1,4) and (5,12)

OpenStudy (anonymous):

I just need to write an equation for these two points

OpenStudy (tkhunny):

We can't do that, here. Please read the Code of Conduct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I read it. Now what do we do?

OpenStudy (tkhunny):

We would be delighted to help you learn how to do it for yourself.

OpenStudy (anonymous):

k

OpenStudy (tkhunny):

We need the slope. Do you know how to calculate the slope, given two points? It typically appears like one of these: \(\dfrac{Rise}{Run} = \dfrac{Change\;in\;y}{Change\;in\;x} = \dfrac{y_{2} - y_{1}}{x_{2} - x_{1}} = \dfrac{\Delta y}{\Delta x}\) Have you seen ANY of those expressions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can we finish this tomorrow I have to go to bed now

OpenStudy (anonymous):

i wrote down those expressions